# arXiv:2602.23144v2[math-ph] 28 Feb 2026

### v2

@@ Abstract
We investigate variational problems in quantum thermodynamics at positive temperature, in which admissible states are constrained by prescribed outcomes of a finite set of measurements. We solve a problem raised by the recent work [21, Section C] and develop a general mathematical setup which allows a broad class of possible regularizations. Employing methods inspired by non-commutative optimal transport, we analyze the dual formulation of the problem, study the existence and characterization of maximizers, and investigate the qualitative behavior of the model in the zero-temperature limit. In the second part, we develop computational algorithms tailored to this class of variational problems and demonstrate the effectiveness of our approach through applications to quantum information tasks, including quantum state tomography and quantum optimal transport.
@@ 1 Introduction
Our work is inspired and generalized several works that recently appear in the literature on quantum termodynamics and regularization on semi-definite programming. Let us first state our framework.
Let $M\in\mathbb{N}$ be a natural number, $\varepsilon>0$ be a positive real number, $\mathcal{H}$ be a finite-dimensional Hilbert space, $H,Q_{0},Q_{1},\dots,Q_{M}\in\mathrm{H}(\mathcal{H})$ be Hermitian matrices and $(q_{0},q_{1},\dots,q_{M})\in\mathbb{R}^{M+1}$ be real numbers. In this paper we are interested in characterize the solution, develop computational algorithms and study the behavoir when $\varepsilon\to 0^{+}$ for the following class of variational problems
@@
In [21], the authors consider the case when $\varphi(z)=z\ln z$ and the variational problem (1.1) corresponds to the von-Neumman regularization of the ground state energy associated with the Hamiltonian $H$. In this case, the solution of that problem at a giving temperature $\varepsilon>0$ is giving by a Gibbs state of an effective Hamiltonian
$\pi_{\varepsilon}=\exp\left(-H_{\mathrm{eff}}\right)\in\operatorname{\rm{Adm}(\mathbf{Q},\mathbf{q})},$
@@
where $H_{\mathrm{eff}}=H-\frac{\sum_{i=0}^{M}\alpha_{i}Q_{i}}{\varepsilon}$.
A motivation for our work is given by [21, Section C], where the authors outlined the potential impact of studying several different types of regularization, beyond the von Neumann entropy. In this contribution, we answer the question raised in the aforementioned paper and develope a general mathematical framework for convex regularizations, which encompass, for instance, both quadratic regularizations and Tsallis entropies.
There are several quantum problems that can be interpreted as semidefinite programming; two special instances of our framework are given by quantum optimal transport and quantum tomography.
The techniques we employ in this work have recently been successfully adopted in the study of quantum optimal transport (QOT). This non-commutative setting admits a duality theory in analogy to classical optimal transport and can be recast using the language of this work by choosing suitable observables and measurements. QOT extends classical optimal transport theory to quantum systems, where non-commutativity has a physical and chemical origin. For classical theory, see the textbooks  [10, 23, 25]. The initial formulations appeared in non-commutative probability [12] and quantum Markov semigroups [7], and were later tied to quantum information [8, 17], statistical mechanics [3, 4] and electronic structure theory [16]. We refer to Section 6.2 for more details.
A special case of the variational problem  (1.1) (with $\varepsilon=0)$ corresponds to the quantum optimal transport problem [4, 13, 14, 19], as shown in [9, Theorem 3.1], or a quantum entropy-regularized optimal transport [5, 16, 18] (with $\varepsilon>0$).
@@
Well-posedness of the problem. Naturally, conditions may need to be imposed on the observables $Q_{i}$ and the outcomes $q_{j}$ in order to have a nonempty set of admissible operators. For example, assuming that $Q_{0},Q_{1},\dots,Q_{j_{0}}$ are a basis for the set of all observables $\{Q_{i}\}_{i=0}^{M}\subset\mathcal{H}(\operatorname{H})$, we show that a necessary condition for $\operatorname{\rm{Adm}(\mathbf{Q},\mathbf{q})}\neq\emptyset$ is given by
@@
Our approach relies on duality theory and the mathematical tools recently developed in [5, 16, 18], see also [22] for a review on the subject. Using this machinery, we obtain several results about our models, which include the proof of an equivalent dual formulation, analysis of the optimizers (both primal and dual) and the relations between them. Finally, we study the behavior of these models (and their optimizers) in the limit as $\varepsilon\to 0$. This yields to a corresponding quantum model at zero temperature, whose properties and minimizers are in detailed analyzed.
Analysis of the regularized problems. Our first result describes the variational problem (1.1) via a suitably (unconstrained when $\varepsilon>0$) dual formulation – see Theorems 4.2 and 4.5 – and states that
@@
$\pi=\psi^{\prime}\left(\frac{\sum_{i=0}^{M}\alpha_{i}Q_{i}-H}{\varepsilon}\right).$ (1.6)
By choosing the von Neumann entropy regularization, we recover a classical Gibbs-state form $\pi=\exp\left(\frac{\sum_{i=0}^{M}\alpha_{i}Q_{i}-H}{\varepsilon}\right)$ with effective Hamiltonian given by $H^{\prime}=H-\sum_{i=0}^{M}\alpha_{i}Q_{i}$.
Convergence results at zero temperature. In a subsequent section, we analyze and discuss the asymptotic behavior of these variational problems and their optimizers in the limit to zero temperature, meaning as $\varepsilon\to 0$. In this framework, the natural notion of convergence to handle minima and minimizers is given by $\Gamma$-convergence (cf. Section 5.1=. We show that, as $\varepsilon\to 0$, both primal and dual formulation converge to the zero temperature (equivalent) problems
@@
where with the latter inequality constraint between operators is intended in the sense of quadratic forms. Analogously to the positive temperature case, we show that, at zero temperature, optimizers are characterized by a complementary slackness conditions, which in this case assumes the form of
@@
Numerical results. The last section contains numerical simulations for the computation of the maximizer of the dual problem at positive temperature, by means of L-BFGS [20]. The stopping criterion is based on attaining a sufficiently small gradient norm of (1.4) within the available computational budget. We specialize our experiments to quantum optimal transport and quantum tomography. We qualitatively analyze convergence behavior for several values of the temperature parameter $\varepsilon$. The numerical results illustrate both the practical applicability of the proposed methods and the computational challenges that may arise.
@@ 2. Case 2:
5 Finite temperature limit ($\varepsilon\to 0$)
In this section, we discuss the behavior of the problems defined in 1.1 and 1.4 when $\varepsilon\to 0$. As usual, we fix $M\in\mathbb{N}$, $\mathcal{H}$ a finite-dimensional Hilbert space, $H,Q_{0},Q_{1},\dots,Q_{M}\in\mathrm{H}(\mathcal{H})$ to be Hermitian matrices over $\mathcal{H}$ and $(q_{0},q_{1},\dots,q_{M})\in\mathbb{R}^{M+1}$. We also pick $Q_{0}\in\operatorname{H}_{>}(\mathcal{H})$ a positive definite Hermitian operator and $q_{0}\in(0,+\infty)$.

### v3

# Quantum thermodynamics and semidefinite programming: regularization and algorithms

@@ Abstract
We investigate variational problems in quantum thermodynamics at positive temperature, in which admissible states are constrained by prescribed outcomes of a finite set of measurements. We solve a problem raised by the recent work [18, Section C] and develop a general mathematical setup which allows a broad class of possible regularizations. Employing methods inspired by non-commutative optimal transport, we analyze the dual formulation of the problem, study the existence and characterization of maximizers, and investigate the qualitative behavior of the model in the zero-temperature limit. In the second part, we develop computational algorithms tailored to this class of variational problems and demonstrate the effectiveness of our approach through applications to quantum information tasks, including quantum state tomography and quantum optimal transport.
@@ 1 Introduction
Let $M\in\mathbb{N}$ be a natural number, $\varepsilon>0$ be a positive real number, $\mathcal{H}$ be a finite-dimensional Hilbert space, $H,Q_{0},Q_{1},\dots,Q_{M}\in\mathrm{H}(\mathcal{H})$ be Hermitian matrices and $(q_{0},q_{1},\dots,q_{M})\in\mathbb{R}^{M+1}$ be real numbers. In this paper we are interested in the following class of variational problems
@@
In [7, 16, 18], the authors consider the case when $\varphi(z)=z\ln z$ and the variational problem (1.1) corresponds to the von Neumman entropy regularization of the ground state energy associated with the Hamiltonian $H$. In this case, the solution of that problem at a giving temperature $\varepsilon>0$ is giving by a Gibbs state of an effective Hamiltonian
@@
$\pi_{\varepsilon}=\exp\left(-\frac{1}{\varepsilon}H_{\mathrm{eff}}\right),\quad\text{where }H_{\mathrm{eff}}=H-\sum_{i=0}^{M}\alpha^{\varepsilon}_{i}Q_{i},\vskip-5.69054pt$ (1.3)
where $(\alpha_{i}^{\varepsilon})_{i=0}^{M}$ are real numbers and can be interpreted as Lagrange multipliers associated with the constraints. In this sense, $H_{\mathrm{eff}}$ plays the role of an effective Hamiltonian, modified by the constraints, such that the corresponding Gibbs state $\pi_{\varepsilon}\in\operatorname{\rm{Adm}(\mathbf{Q},\mathbf{q})}$ belongs to the admissible class of states.
A motivation for our work is provided by [18, Section C], where the authors emphasize the importance of developing a rigorous mathematical theory and computational algorithms capable of handling regularization schemes beyond the von Neumann entropy. In particular, they propose hybrid quantum–classical algorithms to solve the quantum optimization problem (1.1) under von Neumann entropy regularization. While substantial progress has been achieved for Gibbs state-based methods – largely due to their variational characterization and the relative maturity of thermal state preparation techniques – the extension beyond the Gibbs energy/von Neumann entropy regularization framework poses significant conceptual and technological challenges.
In this paper, we address the question raised in the aforementioned work by developing a general mathematical framework for convex regularizations, encompassing, for example, quantum quadratic regularizations and Tsallis-type entropies.
Our approach is grounded in duality theory and builds on mathematical tools that have recently been developed and successfully applied in the study of quantum optimal transport (QOT) problems, a special case of the variational problem (1.1[5, 13, 14, 19] (with $\varepsilon>0)$ and [3, 10, 11, 15] (with $\varepsilon=0)$. We refer to Section 6.2 for more details and also [20] for a review on the subject.
Within this framework, we establish several results for the variational problem (1.1), including: (i) the derivation of an equivalent dual formulation, together with a detailed analysis of both primal and dual optimizers and a clarification of the relations between them; (ii) the investigation of the asymptotic behavior of the problem and its optimizers as $\varepsilon\to 0$, recovering in the limit a corresponding quantum formulation at zero temperature, whose structural properties and minimizers are analyzed.
Finally, we illustrate our methodology by proposing a novel algorithm computing numerical realizations of the dual problem of (1.1) via L-BFGS for both the von Neumman entropy and quadratic quantum entropy regularizations. From our knowledge, (classical) algorithms developed previously focus on the particular case of the von Neumman entropy and are based on first-order methods for dual problem [4, 16, 19, 21].
A promising direction for future development lies in extending hybrid quantum–classical algorithms beyond the von Neumann entropy setting [18] to encompass more general convex regularizations. Achieving this goal would considerably broaden the scope of quantum optimization methods, enabling the treatment of models with quadratic penalties, Tsallis-type entropies, or other non-Gibbsian regularizers that arise naturally in quantum information theory, statistical mechanics, and machine learning.
1.1 Contributions and organization of the paper
For the reader’s convenience, Section 2 presents the complete mathematical framework and the precise statements of the main results of the paper. The remainder of the paper is devoted to their proofs. Below, we provide an overview of our contributions and the organization of this work.
@@ 1.1 Contributions and organization of the paper
Well-posedness of the problem (Section 3). Naturally, conditions may need to be imposed on the observables $Q_{i}$ and the outcomes $q_{j}$ in order to have a nonempty set of admissible operators. For example, assuming that $Q_{0},Q_{1},\dots,Q_{j_{0}}$ are a basis for the set of all observables $\{Q_{i}\}_{i=0}^{M}\subset\mathcal{H}(\operatorname{H})$, we show that a necessary condition for $\operatorname{\rm{Adm}(\mathbf{Q},\mathbf{q})}\neq\emptyset$ is given by
@@
Analysis of the regularized problems (Section 4). Our first result describes the variational problem (1.1) via a suitably (unconstrained when $\varepsilon>0$) dual formulation – see Theorems 4.2 and 4.5 – and states that
@@
$\pi_{\varepsilon}=\psi^{\prime}\left(\frac{\sum_{i=0}^{M}\alpha_{i}Q_{i}-H}{\varepsilon}\right).$ (1.7)
Notice that, by choosing the von Neumann entropy regularization, we recover a Gibbs-state form described in equation (1.3).
Convergence results at zero temperature (Section 5). In a subsequent section, we investigate the asymptotic regime of the above variational problems as the temperature vanishes, that is, in the limit $\varepsilon\to 0^{+}$. This corresponds to the physically relevant zero-temperature limit and leads to a qualitatively different (primal and dual) optimization problems, where entropic effects disappear and only energetic constraints remain.
To rigorously analyze this singular limit, we employ the framework of $\Gamma$-convergence (see Section 5.1). For readers less familiar with this notion, we stress that $\Gamma$-convergence is a variational convergence theory specifically designed to study the stability of minimization problems. Its strength lies in the fact that it guarantees not only the convergence of the functionals themselves, but the convergence of minimizers. In other words, it provides a mathematically robust way to pass to the limit in optimization problems.
Within this framework, we prove that, as $\varepsilon\to 0^{+}$, both the primal and dual formulations converge to the zero-temperature problem
@@
where with the latter inequality constraint between operators is intended in the sense of quadratic forms. This zero-temperature limit problem can be interpreted as a semi-definite programming problem.
We also deduce the convergence (up to subsequences) of minimizers $\Gamma_{\varepsilon}$ toward a minimizer $\Gamma_{0}$ of the zero-temperature problem. This highlights the practical relevance of the theory: the zero-temperature model is not merely a formal limit, but genuinely describes the asymptotic behavior of optimal states of the regularized problem.
Finally, analogously to the positive temperature case, we show that, at zero temperature, optimizers are characterized by a complementary slackness conditions, which in this case assumes the form of
@@
Numerical results (Section 6). The last section contains numerical simulations for the computation of the maximizer of the dual problem at positive temperature, by means of L-BFGS [17]. We specialize our experiments to quantum optimal transport and quantum tomography. We qualitatively analyze convergence behavior for several values of the temperature parameter $\varepsilon$. The numerical results illustrate both the practical applicability of the proposed methods and the computational challenges that may arise.
@@ 2. Case 2:
5 Finite temperature limit ($\varepsilon\to 0^{+}$)
In this section, we discuss the behavior of the problems defined in 1.1 and 1.5 when $\varepsilon\to 0$. As usual, we fix $M\in\mathbb{N}$, $\mathcal{H}$ a finite-dimensional Hilbert space, $H,Q_{0},Q_{1},\dots,Q_{M}\in\mathrm{H}(\mathcal{H})$ to be Hermitian matrices over $\mathcal{H}$ and $(q_{0},q_{1},\dots,q_{M})\in\mathbb{R}^{M+1}$. We also pick $Q_{0}\in\operatorname{H}_{>}(\mathcal{H})$ a positive definite Hermitian operator and $q_{0}\in(0,+\infty)$.